# Spectral Properties of Irradiated Circumbinary Disks around Binary Black Holes Governed by Hydrogen Opacities Dependent on Temperature and Density

### v1

@@ 5 Discussion
Therefore, GWs emitted from the BBH systems modeled in this work are unlikely to be detected by current or near-future GW observatories. In future work, we aim to extend our model to explore systems with either higher frequency or stronger GW emission, such as stellar-mass BBHs detectable by LIGO/Virgo/KAGRA or supermassive BBH systems within LISA’s sensitivity range. Such extensions will allow us to construct and analyze physically consistent models that are potentially observable through both GW and EM channels.
6 Conclusions
@@
    The CBD surface temperature profiles remain largely consistent across different opacity models in the regions dominated by viscous heating. However, significant difference occurs in the midplane temperature profiles due to the substantial influence of disk opacity.
@@
Our results underscore the critical importance of accurately modeling opacity in astrophysical disk systems, highlighting the necessity for incorporating more realistic opacity prescriptions in future theoretical studies. This advancement will significantly enhance our ability to interpret electromagnetic signals associated with gravitational wave events from binary black hole mergers.
@@ Acknowledgments
S.B. acknowledges support from the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (grant RS-2024-00460704). K.H. and S.B. were also supported by the same program (grant 2020R1A2C1007219). This work was conducted during the research year of Chungbuk National University in 2021. This work was partially supported by NSF grant PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), and by the Grant-in-Aid for Scientific Research from MEXT/JSPS of Japan (grant JP21K03619 to A.T.O.).
@@ A.1 Optical depth prescriptions
where $x_{e}=1$ for $\tau_{20}$ and $x_{e}=0$ for $\tau_{31}$ and $\tau_{32}$ are adopted so that
@@ A.2 Analytical solutions
For the terms with the $1$st order of $\beta$, we get the following identities:

### v2

@@ V.2 Orbital Decay, Circumbinary Disk Decoupling, and GW Detection Across Mass Scales
Let us briefly discuss dynamical decoupling for all these mass scales. Following equation (LABEL:eq:adec), the decoupling radius in units of Schwarzschild radius is estimated to be $a_{\rm dec}/r_{\rm S}\sim 70$ for the fixed dimensionless parameters ($q$, $C_{\rm gap}$, $\alpha_{\rm SS}$, and $Y_{\rm in}$) and thus is independent of the binary mass $M$ (although $a_{\rm dec}\propto M$ in physical units). For our fiducial parameters, both $a=100\,r_{\rm S}$ and $a=1000\,r_{\rm S}$ lie in the pre-decoupling regime, even when the system is scaled to SMBH binaries. Once the binary shrinks to $a\lesssim a_{\rm dec}$, the inflow into the central cavity is suppressed, the structure and luminosity of any minidisks around each black hole change significantly, and post-merger inflow of the CBD toward the remnant black hole may produce an afterglow-like rebrightening (Milosavljević & Phinney, 2005). EM emission associated with the LIGO band and the high-frequency end of the LISA band therefore requires a different class of time-dependent disk models that explicitly include CBD decoupling and evolving accretion (e.g., Armitage & Natarajan 2002; Dittmann et al. 2023), whereas the framework developed in this work is mainly designed to describe the earlier, strongly coupled phase.
V.3 Light-curve Periodicity, Time-averaged SED, and Disk Structure Assumptions
Early analytic work on CBD-assisted binary evolution and its electromagnetic signatures (e.g., Armitage & Natarajan 2002, 2005; Milosavljević & Phinney 2005; Haiman et al. 2009; Hayasaki 2009) motivated hydrodynamic studies of circumbinary accretion, which established that streams penetrating the central cavity can feed minidisks in a strongly time-dependent manner (e.g., Hayasaki et al. 2007, 2008; MacFadyen & Milosavljević 2008; Cuadra et al. 2009). Self-consistent azimuthally averaged steady-state solutions for strongly perturbed CBDs were later derived analytically (Kocsis et al., 2012a, b). Subsequent 2D/3D hydrodynamic and MHD simulations and parameter surveys have clarified that the variability is often concentrated near characteristic periods of $\sim 5\,P_{\rm orb}$ for near-circular equal-mass binaries and $\sim P_{\rm orb}$ for eccentric systems, with additional beat-frequency components and pronounced minidisk asymmetries depending on mass ratio, disk thickness, thermodynamics, viscosity, and magnetic flux (e.g., D’Orazio et al. 2013; Shi et al. 2012; Shi & Krolik 2015; Muñoz & Lai 2016; Miranda et al. 2017; Moody et al. 2019; Duffell et al. 2020; Bowen et al. 2018, 2019; Dittmann & Ryan 2022; Most & Wang 2024; Lai & Muñoz 2023). Our focus, however, is the time-averaged SED obtained from this 1D, steady, orbit-averaged CBD model. At radii relevant for reprocessing (e.g., $r\sim 10\,a$), the thermal timescale $t_{\rm th}\simeq(\alpha\Omega)^{-1}$ is $\gg P_{\rm orb}$, and as a result, short-period fluctuations are strongly smoothed. Hence, binary imprints appear primarily through irradiation geometry and inner truncation, shaping the global SED rather than producing sharp periodic features. A fully time-dependent treatment that follows modulated accretion rates and cavity evolution is left for future work.
In this work, we have adopted the assumption that the two black holes have equal masses so that our prescription (15) implies equal accretion rates onto the two minidisks. However, recent parameter-space surveys show in detail how accretion-rate splitting depends on the mass ratio, disk thickness, and viscosity, providing guidance beyond the equal-splitting assumption (e.g., Dittmann & Ryan 2024). These prescriptions and their variants are already employed in population and detectability studies of binary SMBHs (Kelley et al., 2019). Future studies will address the incorporation of these parameter dependencies into our framework.
Finally, we note that the structural equations employ gas pressure only (equations (12) and (17)–(20)); radiation and magnetic pressure are omitted. For hotter or near-Eddington disks where radiation pressure can dominate, the vertical structure and $T_{\rm s}$ may change. Extending the framework to include radiation pressure is a natural next step.
VI Conclusions
@@
    The CBD surface temperature profiles remain largely consistent across different opacity models in the regions dominated by viscous heating. However, significant differences occur in the midplane temperature profiles due to the substantial influence of disk opacity.
@@
  6. 6.
    The CBD spectral shape is primarily controlled by $\xi_{\rm out}$, $M$, and $a$. In particular, $\xi_{\rm out}$ sets the emitting area of the irradiation-powered IR bump and thus its low-frequency extent, while $M$ and $a_{0}$ largely determine the global and inner-disk frequency scales, respectively.
  7. 7.
    For binaries at $a=1000\,r_{\rm S}$, the GW signal lies in the LISA band and the detectability depends strongly on mass: the C2 ($M=100\,M_{\odot}$) and C3 ($M=1000\,M_{\odot}$) cases are detectable with high SNRs at $D_{\rm L}\simeq 10\,{\rm Mpc}$, whereas the C1 ($M=10\,M_{\odot}$) case remains sub-threshold except for very nearby events. This favors joint EM–GW follow-up for intermediate-mass systems.
  8. 8.
    When the irradiation-heated CBD scalings are extrapolated to supermassive BBH masses, the outer surface temperature decreases to $T_{\rm s}\lesssim{\rm a\ few}\times 10^{2}\,{\rm K}$, where dust and metal opacities dominate. Quantitative predictions for the CBDs therefore require composition-dependent tabulated Rosseland-mean opacities beyond the hydrogen-only framework adopted here.
Our results underscore the critical importance of accurately modeling opacity in astrophysical disk systems, highlighting the necessity for incorporating more realistic opacity prescriptions in future theoretical studies. This advancement will significantly enhance our ability to interpret EM signals associated with GW events from BBHs approaching coalescence.
@@ Acknowledgments
We thank the referee for the constructive and generous suggestions that have improved the paper. S.B. acknowledges support from the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (grant RS-2024-00460704). This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (2020R1A2C1007219 and RS-2025-23323627).
This work was also conducted during the research year of Chungbuk National University in 2021. This work was partially supported by NSF grant PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), and by the Grant-in-Aid for Scientific Research from MEXT/JSPS of Japan (grant JP21K03619 to A.T.O.).
@@ A.1 Optical depth prescriptions
where $x_{e}=1$ for $\tau_{20}$ and $x_{e}=0$ for $\tau_{30}$ and $\tau_{31}$ are adopted so that
@@ A.2 Analytical solutions
For the terms with the 11st order of $\beta$, we get the following identities: